Find an equation of the parabola that has the indicated vertex and whose graph passes through the given point
Vertex: (2, -1); point: (4, -3)

Answer by FrankM(1040)

Vertex form -
y-k=a(x-h)^2
enter the given vertex -
y+1=a(x-2)^2 (<<< you see why it's y+1. y- (-1) is y+1)
Now we have the equation with one unknown to solve, "a". So use the given point.
-3+1=a(4-2)^2
-2=a *4
a= -1/2
y+1=-1/2(x-2)^2
A graph to help illustrate the results